OpenStudy (theoreo):

I'm looking for the intersection points of a pair of polar curves r=cos(3*theta) and r= sin(3*theta) on the interval of 0 <= theta <= pi. I know that to find the points I set the equations equal to one another. So I did that and got tan(3*theta) = 1 and I got that theta = pi/12. There are 2 other intersection points that I need to find, but I forgot the method for solving for multiple values. Can somebody give me a refresher?

OpenStudy (freckles):

when is cos(u)=sin(u)?

OpenStudy (freckles):

think about the unit circle

OpenStudy (theoreo):

When tan(u) = 1.... So at 45 degree angles?

OpenStudy (freckles):

well there are other numbers where cos and sin values are the same

OpenStudy (freckles):

so if we replace 3 theta with u then theta=u/3 so that means 0<=u/3<=pi this means you need to solve cos(u)=sin(u) on the interval u in [0,3pi]

OpenStudy (freckles):

so you should get 3 answers

OpenStudy (freckles):

you already have one which is at pi/4

OpenStudy (freckles):

u=pi/4 you could go ahead and say u=pi/4+2pi but there is one more look in the quadrant where both sin and cos are negative

OpenStudy (theoreo):

Ahh I see what you're saying. Ok, so there's pi/4, 5pi/4, and 9pi/4?

OpenStudy (freckles):

\[u=\frac{\pi}{4} \\ u=\frac{5\pi}{4} \\ u=\frac{9 \pi}{4} \\ \text{ so } u=3 \theta \\ 3 \theta= \frac{\pi}{4} \\ 3 \theta=\frac{5 \pi}{4} \\ 3 \theta=\frac{ 9 \pi}{4}\]

OpenStudy (freckles):

solve those three equations

OpenStudy (theoreo):

pi/12, 5pi/12, and 9pi/12. So the method for this or any problem would be to set your theta with coefficient value equal to u, solve for theta, and then substitute that into the boundary inequality?

OpenStudy (freckles):

yeah when i have something like sin(6 x)=1/2 and i want to solve it on the interval 0<x<9 pi I replace 6x with u so so x=u/6 replace x in the inequality above to find out what u needs to be between and then solve sin(u)=1/2 on 0<x<54 pi which this question would be ridiculous because that domain is huge

OpenStudy (freckles):

\[u=\frac{\pi}{6}+2 \pi n \\ u=\frac{5 \pi}{6}+ 2 \pi n \] where n=0,1,...,27

OpenStudy (freckles):

then replace u with 6x and then solve for x

OpenStudy (theoreo):

Ok great, thanks a lot Freckles

OpenStudy (freckles):

actually I think that should be n=0,..,26

OpenStudy (freckles):

but yeah I think you get the idea